# QED corrections of orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$ for HD+ rovibrational transitions beyond Born-Oppenheimer approximation

### 09 Mar 2026 v1

@@ III.2 Singularity in recoil correction
$S_{\text{rec}}$ will be regularized as $R_{\text{rec}}$, namely regular part of $S_{\text{sec}}$, by eliminating singularity of high-energy contribution $H_{H}^{(6)}$. Subsequently, the regularization of recoil contribution will rise a term determined by the $\delta$-function distribution. In the cut-off regularization, it is expressed as
$H_{\text{reg}}=\sum_{a}C_{r}\frac{z_{a}^{3}}{m_{a}}\pi\delta(\mathbf{r}_{a})\,.$ (50)
@@
${E^{\prime}}^{(6)}_{\text{rel-rec}}=\langle\delta H^{\prime}_{4}+\delta H^{\prime}_{5}+\delta H^{\prime}_{6}+R_{\text{rec}}+H_{\text{reg}}+{H^{\prime}}^{\text{1st}}_{R}\rangle\,$, (51)
@@
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ $=\sum_{a,b\neq a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Vp_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle\,.$ (55)
It should be noted that the regular part $R_{\text{rec}}$ of the divergent matrix elements appeared in $S_{\text{rec}}$ (III.2) should be retained after the regularization.
@@
We now sum all contributions from Eqs. (16), (19), (20), (24), (32), (50) and (51) to obtain the total corrections at orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$:
@@
Comparing with previous $E^{(6)}$ of Eq. (25), the relativistic correction due the effective Hamiltonian $H_{\text{rel}}^{(6)}$ has been regularized and split into nonrecoil correction $E_{\text{rel}}^{(6)}$ (32) and recoil correction $E_{\text{rel-rec}}^{(6)}$ (51). It should be noted that the high-energy relativistic correction, so called pure recoil correction $H_{H}^{(6)}$, has been embraced in the contribution arising in the regularization $H_{\text{reg}}$ (50). On the other hand, the second-order corrections ${E}_{\text{sec}}^{(6)}$ is regularized as ${E^{\prime}}_{\text{sec}}^{(6)}=E^{\prime}_{B}+E_{S}+E^{\prime}_{R}$ where $E^{\prime}_{B}=\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH^{\prime}_{B}\rangle$, $E^{\prime}_{R}=2\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH_{R}\rangle$ and $E_{S}$ (24) are regular enough for numerical study.
@@ IV Numerical Calculations and Conclusion
(0,0) (0,1) (1,0) (1,1)
$\langle\delta H^{\prime}_{1}\rangle$ –1.8703 480(6) –1.8137 755(5) –1.8687 273(6) –1.8122 633(5)
$\langle\delta H^{\prime}_{2}\rangle$ 1.7125 566(5) 1.6630 255(5) 1.7109 929(5) 1.6615 589(5)
$\langle\delta H^{\prime}_{3}\rangle$ –0.0000 0628 5744(2) –0.0000 0603 2892(2) –0.0000 0626 5720(2) –0.0000 0601 3395(2)
$\langle\delta{H^{\prime}}_{B}^{\text{1st}}\rangle$ 0.4020 166(1) 0.3880 342(1) 0.4016 656(1) 0.3877 119(1)
${E^{\prime}}_{\text{rel}}^{(6)}$ 0.2442 189(8) 0.2372 782(7) 0.2439 249(8) 0.2370 001(7)
$\langle\delta H^{\prime}_{4}\rangle$ 0.0007 150(4) 0.0007 015(4) 0.0007 144(4) 0.0007 010(4)
$\langle\delta H^{\prime}_{5}\rangle$ 0.0005 5890 9980 0.0005 4532 3594 0.0005 5831 3896 0.0005 4475 9030
$\langle\delta H^{\prime}_{6}\rangle$ –0.0002 1032 7438 –0.0002 0656 2370 –0.0002 1021 536(2) –0.0002 0645 96(2)
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ –0.0002 1544 0395 –0.0002 1617 5234(2) –0.0002 1533 761(1) –0.0002 1606 7579(5)
$\langle R_{\text{rec}}\rangle$ 0.0000 3966 5983 0.0000 3761 7674 0.0000 4008 5121 0.0000 3757 6478
$\langle H_{\text{reg}}\rangle$ 0.0004 1079 5750 0.0004 0137 2924 0.0004 1042 9068 0.0004 0102 6584
${E^{\prime}}_{\text{rel-rec}}^{(6)}$ 0.0012 986(4) 0.0012 630(4) 0.0017 977(4) 0.0012 618(4)
@@
The $m\alpha^{6}$ order corrections to transition frequencies in the HD+ ion are presented in Table 4. The uncertainty in the nonrecoil relativistic correction $\Delta{E^{\prime}}_{\text{rel}}^{(6)}$ and the recoil relativistic correction $\Delta{E^{\prime}}_{\text{rel-rec}}^{(6)}$ are estimated to be 20 Hz and 11 Hz respectively, primarily due to neglect of the higher-order recoil effects. The uncertainty due to the second-order contribution $\Delta(E^{\prime}_{B}+E_{S})$ is about 26 Hz, which is estimated by the last significant digit of the corresponding values of $E^{\prime}_{B}$ and $E_{S}$. The resulting $m\alpha^{6}$-order correction to the $(0,0)\!\to\!(0,1)$ transition is determined to be $1710.684(35)~\text{kHz}$, the uncertainty of which is three orders of magnitude smaller than the value reported in previous theoretical studies [21, 23]. As is seen, there is a discrepancy of approximately 1.8 kHz between the obtained value and the previous theory. We attribute this to the limitations of the adiabatic Born-Oppenheimer approximation employed in previous works.
@@ Appendix C Derivation of modified effective Hamiltonian
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle=\sum_{b\neq a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Vp_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle+O(1/m_{p,d}^{2})\,$, (155)

### 16 Mar 2026 v2

@@ III.1 Singularity in nonrecoil correction
$\langle\delta H^{\prime}_{1}\rangle$ $=\frac{1}{16m_{e}^{4}}\bigg\{2E_{0}\langle p_{e}^{4}\rangle-\langle\{p_{e}^{4},V_{12}\}\rangle-\sum_{a}\bigg[\langle\{p_{e}^{4},V_{a}\}\rangle^{\prime}$
$+\frac{4m_{e}^{2}}{m_{a}}\left(E_{0}^{2}\langle p_{a}^{2}\rangle-2E_{0}\langle Vp_{a}^{2}\rangle+\langle Vp_{a}^{2}V\rangle^{\prime}\right)\bigg]\bigg\}\,$, (38)
$\langle\delta H^{\prime}_{2}\rangle$ $=\sum_{a}\bigg[-\frac{1}{32m_{e}^{3}}\bigg(\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle\mp\frac{m_{e}}{m_{a}}\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle\bigg)$
$+\frac{3}{64m_{e}^{4}}\langle\{p_{e}^{2},\rho_{a}\}\rangle^{\prime}\bigg]\,$, (39)
$\langle\delta H^{\prime}_{3}\rangle$ $=\sum_{a}\frac{1}{8m_{e}^{2}m_{a}}(\mp)\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle\,$, (40)
@@ III.2 Singularity in recoil correction
$S_{\text{rec}}$ will be regularized as ${(S_{\text{rec}})}_{r}$, namely regular part of $S_{\text{sec}}$,
${(S_{\text{rec}})}_{r}=\sum_{a}\frac{1}{m_{e}^{2}m_{a}}\left(\frac{\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}}{8}-\frac{m_{e}\langle V_{a}^{3}\rangle_{r}}{4}-\frac{\langle V_{a}\rho_{a}\rangle_{r}}{8}\right)\,.$ (53)
by eliminating singularity of high-energy contribution $H_{H}^{(6)}$. Subsequently, the regularization of recoil contribution will rise a term determined by the $\delta$-function distribution. In the cut-off regularization, it is expressed as
$H_{r}=\sum_{a}C_{r}\frac{z_{a}^{3}}{m_{a}}\pi\delta(\mathbf{r}_{a})\,.$ (54)
@@
${E^{\prime}}^{(6)}_{\text{rel-rec}}$ $=\langle\delta H^{\prime}_{4}+\delta H^{\prime}_{5}+\delta H^{\prime}_{6}+H_{r}+{H^{\prime}}^{\text{1st}}_{R}\rangle$
$\quad+{(S_{\text{rec}})}_{r}\,$, (55)
@@
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ $=\sum_{a,b\neq a}-\frac{2z_{e}z_{a}}{m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Up_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle\,.$ (59)
It should be noted that the regular part ${(S_{\text{rec}})}_{r}$ of the divergent matrix elements appeared in $S_{\text{rec}}$ (III.2) should be retained after the regularization.
@@
We now sum all contributions from Eqs. (16), (19), (20), (24), (32), (54) and (III.2) to obtain the total corrections at orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$:
@@
Comparing with previous $E^{(6)}$ of Eq. (25), the relativistic correction due the effective Hamiltonian $H_{\text{rel}}^{(6)}$ has been regularized and split into nonrecoil correction $E_{\text{rel}}^{(6)}$ (32) and recoil correction $E_{\text{rel-rec}}^{(6)}$ (III.2). It should be noted that the high-energy relativistic correction, so called pure recoil correction $H_{H}^{(6)}$, has been embraced in the contribution arising in the regularization $H_{r}$ (54). On the other hand, the second-order corrections ${E}_{\text{sec}}^{(6)}$ is regularized as ${E^{\prime}}_{\text{sec}}^{(6)}=E^{\prime}_{B}+E_{S}+E^{\prime}_{R}$ where $E^{\prime}_{B}=\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH^{\prime}_{B}\rangle$, $E^{\prime}_{R}=2\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH_{R}\rangle$ and $E_{S}$ (24) are regular enough for numerical study.
@@ IV Numerical Calculations and Conclusion
$\langle z_{e}z_{1}p_{e}^{i}\mathcal{W}_{1}^{ij}Up_{1}^{j}\rangle$ 0.3795 4851 4771 0.3700 4007 8771 0.3791 6464 1365 0.3696 7653 009(1)
$\langle z_{e}z_{2}p_{e}^{i}\mathcal{W}_{2}^{ij}Up_{2}^{j}\rangle$ 0.3784 3817 9175 0.3682 5701 6903 0.3780 2354 8783(1) 0.3678 6502 5691(6)
@@
Eqs. (0,0) (0,1) (1,0) (1,1)
$\langle\delta H^{\prime}_{1}\rangle$ (III.1) –1.8703 480(6) –1.8137 755(5) –1.8687 273(6) –1.8122 633(5)
$\langle\delta H^{\prime}_{2}\rangle$ (39) 1.7125 566(5) 1.6630 255(5) 1.7109 929(5) 1.6615 589(5)
$\langle\delta H^{\prime}_{3}\rangle$ (40) –0.0000 0628 5744(2) –0.0000 0603 2892(2) –0.0000 0626 5720(2) –0.0000 0601 3395(2)
$\langle\delta{H^{\prime}}_{B}^{\text{1st}}\rangle$ (43) 0.4020 166(1) 0.3880 342(1) 0.4016 656(1) 0.3877 119(1)
${E^{\prime}}_{\text{rel}}^{(6)}$ (32) 0.2442 189(8) 0.2372 782(7) 0.2439 249(8) 0.2370 001(7)
$\langle\delta H^{\prime}_{4}\rangle$ (56) 0.0007 150(4) 0.0007 015(4) 0.0007 144(4) 0.0007 010(4)
$\langle\delta H^{\prime}_{5}\rangle$ (57) 0.0005 5890 9980 0.0005 4532 3594 0.0005 5831 3896 0.0005 4475 9030
$\langle\delta H^{\prime}_{6}\rangle$ (58) –0.0002 1032 7438 –0.0002 0656 2370 –0.0002 1021 536(2) –0.0002 0645 96(2)
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ (59) –0.0002 1544 0395 –0.0002 1617 5234(2) –0.0002 1533 761(1) –0.0002 1606 7579(5)
${(S_{\text{rec}})}_{r}$ (53) 0.0000 3966 5983 0.0000 3761 7674 0.0000 4008 5121 0.0000 3757 6478
$\langle H_{r}\rangle$ (54) 0.0004 1079 5750 0.0004 0137 2924 0.0004 1042 9068 0.0004 0102 6584
${E^{\prime}}_{\text{rel-rec}}^{(6)}$ (III.2) 0.0012 986(4) 0.0012 630(4) 0.0017 977(4) 0.0012 618(4)
@@
The $m\alpha^{6}$ order corrections to transition frequencies in the HD+ ion are presented in Table 4. The uncertainty in the nonrecoil relativistic correction $\Delta{E^{\prime}}_{\text{rel}}^{(6)}$ and the recoil relativistic correction $\Delta{E^{\prime}}_{\text{rel-rec}}^{(6)}$ are estimated to be 20 Hz and 11 Hz respectively, primarily due to neglect of the higher-order recoil effects. The uncertainty due to the second-order contribution $\Delta(E^{\prime}_{B}+E_{S})$ is about 26 Hz, which is estimated by the last significant digit of the corresponding values of $E^{\prime}_{B}$ and $E_{S}$. The resulting $m\alpha^{6}$-order correction to the $(0,0)\!\to\!(0,1)$ transition is determined to be $1710.684(35)~\text{kHz}$, the uncertainty of which is three orders of magnitude smaller than the value reported in previous theoretical studies [21, 23]. As is seen, there is a discrepancy of approximately 1.8 kHz between the obtained values and the previous theory on transitions $(0,0)\!\to\!(0,1)$ and $(0,0)\!\to\!(1,1)$ as shown in Table 4. We attribute this to the limitations of the adiabatic Born-Oppenheimer approximation employed in previous works.
@@ Appendix C Derivation of modified effective Hamiltonian
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ $=\sum_{b\neq a}-\frac{2z_{e}z_{a}}{m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Up_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle$
$\quad+O(1/m_{p,d}^{2})\,$, (159)